# Matthias von Davier

### v2

@@ 2.1.4 Improvement for Small $\nu_{k}$ and Small $K$
The following tables show results for sums of $K$ simple variance estimates $S_{k}^{2}$ with
$\frac{\nu_{k}S_{k}^{2}}{\sigma_{k}^{2}}$
$\chi^{2}$ distributed. The ideal case where $\forall k,j:\sigma_{k}^{2}=\sigma_{j}^{2},\nu_{k}=\nu_{j},w_{k}=w_{j}$ was used to be able to compare the original Satterthwaite formula to the improved estimator, and against the expected value
@@
that follows in this case since each component $S_{k}^{2}$ has the same distribution.
@@
The obvious conclusion is that the correction of the Satterthwaite effective sample size is necessary even for large component degrees of freedom. The case where $K=64$ and $\nu_{k}=\overline{\nu}=32$ shows this clearly. The average $df$ for the corrected estimate is $\overline{\nu_{corr}^{*}}=2048.5\approx 2028=K\times\overline{\nu}>\overline{\nu_{satt}^{*}}=1929.9$. The other end of the table shows for $K=2$ and $df=1$ that $\overline{\nu_{corr}^{*}}=2.23\approx 2=K\times\overline{\nu}>1.41=\overline{\nu_{satt}^{*}}.$
@@ 3 Kish’s (1965) Effective Sample Size Estimate
Across the literature, there is another expression that is often given for the effective sample size. This expression is based on the relvariance, which stands for relative variance, of the weights. It is the variance of the weights divided by the average of the weights. That is
$D_{eff}=\left[1+cv^{2}\left(\boldsymbol{w}\right)\right]$
@@ 5 Examples Beyond Effective Sample Size
5.1 Jackknifing and Balanced Repeated Replications (BRR)
In the case of single component d.f. $\nu_{k}=1$ and large(-ish) $K$ representing the (non-trivial) jackknife zones we have no weights to account for and obtain
$\nu^{*}\approx\frac{3\left(\sum_{k=1}^{K}\left[T_{\cdot}-T_{k}\right]^{2}\right)^{2}}{\left[T_{\cdot}-T_{k}\right]^{4}}-2$
since $\nu_{k}+2=3$ for all $k.$ Here, the $T_{k}$ are the pseudo-values calculated under the jackknife scheme of dropping units, and $T_{\cdot}=\sum_{k}T_{k}$. In the case of the BRR the $T_{k}$ are the balanced replicated estimates (of the half samples) obtained by adjusting the weights of the pairs and recalculating the statistics for each zone $K$.
@@ 5.2 Total Variance under Multiple Imputations
The total variance of an estimator when imputations is a prime example of a weighted variance estimate. The total variance as defined by Rubin (1987) is given by
@@
where the sampling variance may be estimated according to some resampling scheme (Johnson & Rust, 1992) and the imputation variance is the sample standard deviation across $M$ imputation based calculations of the same statistic.
In this case, the following weights are used:
@@
and the corrected equation can be directly applied.

### v4

# arXiv:2602.20912v4[stat.AP] 02 Mar 2026

@@ 2.1.4 Improvement for Small $\nu_{k}$ and Small $K$
Interestingly, Boardman (1974) discusses a similar estimator for $K=2$,
$\nu^{*}\approx\frac{\left(a_{1}S_{1}^{2}+a_{2}S_{2}^{2}\right)^{2}}{\frac{a_{1}^{2}S_{1}^{4}}{\nu_{1}+2}=\frac{a_{2}^{2}S_{2}^{4}}{\nu_{2}+2}}$
which he refers to as the ’unbiased’ Satterthwaite, but does not appear to derive the result. However, he mentions that Satterthwaite (1941) was aware that his formula is biased for small $\nu_{k}$. Note that Boardman’s (1974) estimator is missing the final additive term 2-2 that corrects the estimator.
@@
The following tables show results for sums of $K$ simple variance estimates $S_{k}^{2}$ with $\frac{\nu_{k}S_{k}^{2}}{\sigma_{k}^{2}}=X_{k,\nu_{k}}^{2}\sim\chi_{\nu_{k}}^{2}.$ The ideal case where $\forall k,j:\sigma_{k}^{2}=\sigma_{j}^{2}=\sigma^{2},\nu_{k}=\nu_{j}=\overline{\nu},w_{k}=w_{j}=\frac{1}{K}$ so that $Kw_{i}=1$ was used to enable comparisons of the original Satterthwaite equation to the improved estimator, and against the expected value
@@
that follows in this case since $\frac{K\overline{\nu}}{\sigma^{2}}S^{2}=\sum_{k=1}^{K}Kw_{i}\frac{\overline{\nu}}{\sigma^{2}}S_{k}^{2}=\sum_{k=1}^{K}X_{k,\overline{\nu}}^{2}\sim\chi_{K\overline{\nu}}^{2}.$
@@
The obvious conclusion is that the correction of the Satterthwaite equation for estimating effective sample size is necessary even for reasonably large component degrees of freedom. The case where $K=64$ and $\nu_{k}=\overline{\nu}=32$ shows this clearly. The average $df$ for the corrected estimate is $\overline{\nu_{corr}^{*}}=2048.5\approx 2028=K\times\overline{\nu}>\overline{\nu_{satt}^{*}}=1929.9$. The other end of the table shows for $K=2$ and $df=1$ that $\overline{\nu_{corr}^{*}}=2.23\approx 2=K\times\overline{\nu}>1.41=\overline{\nu_{satt}^{*}}.$
@@ 3 Kish’s (1965) Effective Sample Size Estimate
Across the literature, there is another expression that is often given for the effective sample size. This expression is based on the relvariance of the weights. That is
$D_{eff}=1+cv^{2}\left(\boldsymbol{w}\right)$
@@ 5 Examples Beyond Effective Sample Size
5.1 Jackknifing
Let $T=T\left(X_{1},...,X_{L}\right)$ be a statistic that is computed based on l=1,,Ll=1,...,L samples $X_{l}$. The jackknife estimate (e.g. Efron & Stein, 1981) of the variance $Var\left(T\right)$ is based on calculating so-called pseudo-values
$T_{k}=T\left(X_{1},\dots,X_{k-1},X_{k+1},\dots,X_{K}\right)$
which compute the same statistic under omission of the observation $X_{k}$. Then, the variance of the statistic $T$ can be estimated as
$Var\left(T\left(X_{1},\dots,X_{K}\right)\right)=\frac{K-1}{K}\sum_{k=1}^{K}\left[T_{k}-T_{\cdot}\right]^{2}$
with $T_{\cdot}=\frac{1}{K}\sum_{k}T_{k}$
Each of the terms represents a variance component $S_{k}^{2}=\frac{K-1}{K}\left[T_{k}-T_{\cdot}\right]^{2}$and $Var\left(T\right)=S^{2}=\sum S_{k}^{2}$ . The constant weight factor $\frac{K-1}{K}$ is eliminated when forming the Satterthwaite expression, and we can assume $\nu_{k}=1$ for all $k$. Hence, we obtain
$\nu^{*}\approx\frac{3\left(\sum_{k=1}^{K}\left[T_{k}-T_{\cdot}\right]^{2}\right)^{2}}{\left[T_{k}-T_{\cdot}\right]^{4}}-2$
since $\nu_{k}+2=1+2=3$ for all $k.$
@@ 5.2 Total Variance under Multiple Imputations
The total variance of an estimator when imputations are used is a prime example of a weighted variance estimate. The total variance is defined as the combination of sampling and imputation variance, and is given by
@@
according to Rubin (1987). Rubin & Schenker (1986) discuss this case and propose an approximation of the effective d.f. by assuming the sampling variance has infinite d.f., while the imputation variance has finite d.f. However, the sampling variance may be estimated according to some resampling scheme (Johnson & Rust, 1992) and the imputation variance is the sample variance across $M$ imputation based calculations of the same statistic. Lipsitz et al. (2002) provide an improved approximation that also takes the finite d.f. of the sampling variance into account. In this case, the following weights can be used
@@
and the corrected equation can be directly applied. That is
$\nu_{total}^{*}\approx\frac{\left(Var\left(sampling\right)+\frac{M+1}{M}Var\left(imputation\right)\right)^{2}}{\frac{Var\left(sampling\right)^{2}}{\nu_{sampling}+2}+\left(\frac{M+1}{M}\right)^{2}\frac{Var\left(imputation\right)^{2}}{\nu_{imputation}+2}}-2$
Note that $\nu_{imputation}=M-1$ and that $\nu_{sampling}$ may require estimation using, for example, the corrected Satterthwaite equation applied to the jackknife variance components as illustrated in the previous subsection.